# Progressive Contrastive Learning with Multi-Prototype for Unsupervised Visible-Infrared Person Re-identification

### v1

@@ 4.2 Results and Analysis
4.3 Ablation Study
In this section, we conduct ablation studies on the SYSU-MM01 dataset in both all-search and indoor-search modes to show the effectiveness of each component in our method. The results are shown in Tab. 2.
Baseline Settings. We use PGM Wu and Ye (2023) as our baseline. Although PGM has achieved a promising performance on the USVI-ReID task, the neglect of hard samples hinders its further improvement.
Effectiveness of HPCL. The HPCL is proposed to mine hard samples at the cluster level. As shown in Tab. 2, When adding the HPCL on Baseline, the performance improves a large margin of 5.8% in Rank-1 and 5.1% in mAP, respectively. It shows that hard samples can be effectively mined using hard prototype contrast learning, facilitating the model to learn more discriminative features.
Effectiveness of DPCL. The DPCL is proposed to mine hard samples at the instance level. The results show that Rank-1 accuracy can be improved by 2.8% in Rank-1 and 2.7% in mAP when adding the DPCL on Baseline, which confirms that contrastive learning with dynamic prototype can further mine hard samples.
Effectiveness of PCL. PCL is introduced to smoothly shift the model’s attention from simple samples to hard ones. The results show that Rank-1 accuracy can be improved by about 1% in Rank-1 and mAP compared to adding simultaneously the HPCL and DPCL on the Baseline. This confirms that progressive contrastive learning plays a valuable role in assisting HPCL and DPCL in mining hard samples.
Surprisingly, contrastive learning with both hard and dynamic prototypes significantly exceeds the baseline by a large margin of 8.1% in Rank-1 and 7.0% in mAP. The HPCL and DPCL can complement each other to mine hard samples, which effectively guides the network to learn more robust and discriminative features.
@@ 4.4 Further Analysis
Hyper-parameter analysis. Hyper-parameter $\lambda$ is a weighting parameter to trade-off $L_{HPCL}$ and $L_{DPCL}$. Fig. 3 (a) presents the results under different values of $\lambda$. We can observe that when $\lambda$ is small, i.e., $L_{DPCL}$ contributes more to the model, the performance degrades. However, when $\lambda$ is large, i.e., $L_{HPCL}$ contributes heavily to the model, the model both achieves superior performance. Note that when $\lambda=1$, i.e., the proposed method is trained without DPCL, the performance drops significantly. $\lambda$ is finally set to 0.5 and our method achieves the best performance of 64.4% in Rank-1. Moreover, we also analyze the effect of the number of hard samples at hard prototype. As shown in Fig. 3 (b), we vary the $k$ from 1 to 3 and keep the other hyper-parameters fixed, which shows that PCLMP achieves the best performance when $k=1$.
The ARI metric. Following MMM Shi et al. (2024), we utilize the Adjusted Rand Index (ARI) metric for clustering evaluation. The larger the ARI value, the higher the clustering quality. In Fig. 3 (c), “RGB” and “IR” denote the ARI values of visible and infrared clusterings, which can measure the quality of visible and infrared pseudo-labels. “ALL” means the ARI values of overall clusterings, which can evaluate the reliability of cross-modality correspondences. PCLMP surpasses other methods significantly on all of the mentioned ARI values, which demonstrates our method can effectively mine hard samples to improve clustering quality.
Visualization. As shown in Fig. 4, we visualize the t-SNE map of 10 randomly chosen identities from SYSU-MM01. Compared to the baseline, the distribution of the same identity from the same modality is more compact and the distance of the same identity from different modalities is closer together. Moreover, some hard samples in the baseline are incorrectly clustered, while these hard samples are well clustered in our method, which shows the effectiveness of the proposed PCLMP.
@@ 5 Conclusions
In this paper, we propose a novel method for USVI-ReID called Progressive Contrastive Learning with Multi-Prototype (PCLMP), which learns both shared and diverse information. To be specific, we design Hard Prototype Contrastive Learning to mine distinctive yet significant information and Dynamic Prototype Contrastive Learning to preserve intrinsic diversity features. Furthermore, we introduce a progressive learning strategy to incorporate both HPCL and DPCL into the model. Extensive experiments demonstrate that PCLMP outperforms state-of-the-art USVI-ReID methods.

### v2

@@ 4.4 Further Analysis
Hyper-parameter analysis. Hyper-parameter $\lambda$ is a weighting parameter to trade-off $L_{HPCL}$ and $L_{DPCL}$. Fig. 2 (a) presents the results under different values of $\lambda$. We can observe that when $\lambda$ is small, i.e., $L_{DPCL}$ contributes more to the model, the performance degrades. However, when $\lambda$ is large, i.e., $L_{HPCL}$ contributes heavily to the model, the model both achieves superior performance. Note that when $\lambda=1$, i.e., the proposed method is trained without DPCL, the performance drops significantly. $\lambda$ is finally set to 0.5 and our method achieves the best performance of 64.4% in Rank-1. Moreover, we also analyze the effect of the number of hard samples at hard prototype. As shown in Fig. 2 (b), we vary the $k$ from 1 to 3 and keep the other hyper-parameters fixed, which shows that PCLHD achieves the best performance when $k=1$. Hard samples are distributed in multiple directions, so multiple hard samples cannot be represented by a single prototype. This is why using more hard samples as prototypes leads to a decline in overall performance
The ARI metric. Following MMM [48], we utilize the Adjusted Rand Index (ARI) metric for clustering evaluation. The larger the ARI value, the higher the clustering quality. In Fig. 2 (c), “RGB” and “IR” denote the ARI values of visible and infrared clusterings, which can measure the quality of visible and infrared pseudo-labels. “ALL” means the ARI values of overall clusterings, which can evaluate the reliability of cross-modality correspondences. PCLHD surpasses other methods significantly on all of the mentioned ARI values, which demonstrates our method can effectively mine divergence and variety to improve clustering quality.
Visualization. As shown in Fig. 3, we visualize the t-SNE map of 10 randomly chosen identities from SYSU-MM01. Compared to the baseline, the distribution of the same identity from the same modality is more compact and the distance of the same identity from different modalities is closer together. Moreover, some hard samples in the baseline are incorrectly clustered, while these hard samples are well clustered in our PCLHD, which shows the effectiveness of the proposed PCLHD.
@@ 5 Conclusions
In this paper, we propose a novel method for USVI-ReID called Progressive Contrastive Learning with hard and dynamic prototype (PCLHD), which learns commonality, divergence and variety. To be specific, we design Hard Prototype Contrastive Learning to mine divergent yet significant information and Dynamic Prototype Contrastive Learning to preserve intrinsic variety features. Furthermore, we introduce a progressive learning strategy to incorporate both HPCL and DPCL into the model. Extensive experiments demonstrate that PCLHD outperforms state-of-the-art USVI-ReID methods.